Assume pipelining and 20 packets can be sent without ACK, 100 Mbps link, 10 ms propagation delay, 10,000 bit packet. Calculate sender utilization.
Answer #1

ANSWER:

GIVEN THAT:

CALCULATING THE SENDER UTILIZATION:

1. Transmission delay (Tt) = length / bandwidth

= 10000/(100* 106)

= 0.0001 sec

= 0.1 ms

2. Propagation delay ( Tp) = 10 ms.

And number of packets(N)  at a time given is 20.

Then a is (Tp/Tt) = 100

So, sender utilization / efficiency

= N / ( 1 + 2a)

= 20 / ( 1 + 2*100)

= 20/ 201

= 9.95 % efficiency